Differentiate f(x)=tan^-1(1-x/1+x)
Differentiate x^x
1
Expert's answer
2020-06-28T18:45:54-0400

i)


"\frac{d(arctan\frac{1-x}{1+x})}{dx}=" "\frac{1}{(\frac{1-x}{1+x})^{2}+1}\frac{d}{dx}\frac{1-x}{1+x}"


="\frac{1}{(\frac{1-x}{1+x})^{2}+1}\frac{(-1)(1+x)-(1)(1-x)}{(1+x)^{2}}"

="\frac{1}{(\frac{1-x}{1+x})^{2}+1}\frac{-2}{(1+x)^{2}}"



simplifying we get


d(tan-1(1-x/1+x) = "\frac{-1}{1+x^{2}}"



ii) d(xx)/dx = "\frac{d}{dx}e^{ln(x) x}"

= exln(x) "\frac{d}{dx}xlnx"


="e^{xlnx}(1*lnx+x*\frac{1}{x})"

= "e^{xlnx}(lnx+1)"


"x^{x}(ln x+1)"